# www.entalus.com

### v1

@@ 7.1 Set Transformers
Definition 11 (Inclusive Predicates)
A stream predicate $P\subseteq{{A}^{\omega,{n}}}$ is said to be inclusive if for any convergent sequence $(f_{k})_{k\in\mathbb{N}}$ of streams in ${{A}^{\omega,{n}}}$ the following holds:
$((\forall{k\in\mathbb{N}})\,{f_{k}\in P})$ $\,\Rightarrow\,(\lim_{k\to\infty}f_{k})\in P$
Corollary 5 (Induction Principle for sp.subscriptnormal-$-.{\mathrm{sp}}_{.}$)
Let $P\subseteq{{A}^{\omega,{n}}}$ be an inclusive predicate and $T:{{A}^{\omega,{n}}}\to{\mathrm{Comp}({{{A}^{\omega,{n}}}})}$ contracting; then
$(\forall{Q\subseteq P})\,{{\mathrm{sp}}_{T}(Q)\subseteq P}$  implies $\text{~{}~{}implies~{}~{}}{\mathrm{fix}}({sp_{T}})\subseteq P\mbox{\@.}$ (26)
Proof
For $F_{0}\,\coloneqq\,\emptyset$ and $F_{k+1}={\mathrm{sp}}_{T}F_{k}$ we get, by induction on $k$ and the assumption above, that $F_{k}\subseteq P$ for all $k\in\mathbb{N}$. Since $P$ is inclusive, ${\mathrm{fix}}({{\mathrm{sp}}_{T}})=\lim_{k\to\infty}F_{k}\subseteq P$.
Remark 20
For the adjointness of ${\mathrm{sp}}_{T}$ and ${\mathrm{wp}}_{T}$ we might also use the equivalent assumption $(\forall{Q\subseteq P})\,{Q\subseteq{\mathrm{wp}}_{T}(P)}$ in the induction principle for ${\mathrm{sp}}_{T}$ (see Corollary 5).
7.2 Contractive Maps
An identically empty mapping $T$ can not have fixed points, since $f\notin\emptyset=T(f)$. Also, we restrict the codomain of multivalued mappings to closed sets only, since any closed subset of a Cauchy complete space is Cauchy complete.
Lemma 11
A nondeterministic contractive stream transformer $T$ has a fixpoint if (1) $T$ is not identically empty and (2) $T(f)$ is closed for all streams $f$.
Proof
Since $T$ is not identically empty, there exist $f_{0},f_{1}\in{A}^{\omega}$ such that $f_{1}\in T(f_{0})$. Now, since $T$ is contractive there is $10\leq l<$ 1 and a $f_{2}\in T(f_{1})$ with $d(f_{1},f_{2})\leq l\cdot d(f_{0},f_{1})$. In this way, we recursively construct a sequence $(f_{k})$ such that $f_{k+1}\in T(f_{k})$ and
@@ 7.2 Contractive Maps
$d(f_{k},f_{m})\leq\mathrm{max}(d(f_{k},f_{k+1}),\ldots,d(f_{m-1},f_{m}))\leq l^{k}\mbox{\@,}$ (27)
@@
$d(f_{k},f^{*})$ $\leq l^{k}\mbox{\@.}$ (28)
Since $T$ is a contraction, there is a sequence $(g_{k})_{k\in\mathbb{N}}$ with $d(f_{k+1},g_{k})\leq l\,d(f_{k},f^{*})$. Therefore, using inequality (28):
@@
Now, since $\lim_{k\to\infty}d(f^{*},g_{k})=0$,
$\lim_{k\to\infty}g_{k}=f^{*}\mbox{\@.}$
But $g_{k}\in T(f^{*})$ and all images of $T$ are closed, and therefore the limit $f^{*}$ of the sequence $(g_{k})$ is an element of $T(f^{*})$. Altogether, $f^{*}\in T(f^{*})$.
In general, fixpoints of contractive multivalued maps in the sense of Definition 7 are not unique, but a slightly stronger requirement on contractiveness implies uniqueness of fixpoints.
Definition 12 (Strong Contractions)
@@
Remark 21
If $\emptyset\notin T({{A}^{\omega,{n}}})$ then every strong $l$-contraction also is a (weak) $l$-contraction in the sense of Definition 7. Moreover, a deterministic stream transformer $T$ is $l$-contractive if and only if it is strongly $l$-contractive.
Lemma 12
@@
Consequently, $d(f^{*},g^{*})=0$, and any two fixpoints of $T$ are equal.
Remark 22
For deterministic maps, the fixpoint iteration with $f_{k+1}\in T(f_{k})$, as constructed in the proof of Lemma 11 reduces to the Picard iteration $f_{k+1}=T(f_{k})$. Moreover, uniqueness of fixpoints for deterministic, $\delta>0$-causal maps can also be shown directly, since the equality of two arbitrary fixpoints $f^{*}$, $g^{*}$ directly follows from $d(f^{*},g^{*})=d(T(f^{*}),T(g^{*}))\leq 2^{-\delta}\cdot d(f^{*},g^{*})$.
The following result directly follows from Lemmata 511, and 12.
Theorem 7.2
A strongly causal stream transformer $T$ has a fixpoint if it is (1) not identically empty, and (2) $T(f)$ is closed for each stream $f$. If, in addition, $T$ is strongly contracting then this fixpoint is unique.
As an immediate consequence of Theorem 7.2 we get an automata-free construction of unique fixpoints for strongly causal deterministic stream transformers over streams with infinite value sets.
Corollary 6
@@
7.3 Fixpoint Induction
The fixpoint iteration for strongly causal stream transformers in Theorem 7.2 is used to derive a fixpoint induction principle. The overall approach is analogous to deriving fixpoint induction for Scott-continuous functions on complete partial orders [62]. for strongly causal stream transformers.
Lemma 13 (Fixpoint Induction)
Let $T$ be a stream transformer as in Theorem 7.2 with a unique fixpoint ${\mathrm{f}ix}(T)$ and $P$ an inclusive stream predicate; then:
$((\forall{f\in P})\,{T(f)\in P})\,\Rightarrow\,{\mathrm{f}ix}(T)\in P\mbox{\@.}$ (29)
Proof
Let $(f_{k})$ be the sequence with $f_{k+1}\in T(f_{k})$ for $k\in\mathbb{N}$ and ${\mathrm{f}ix}(T)=\lim_{k\to\infty}f_{k}$ as constructed in the proof of Lemma 11. From the assumption and natural induction on $k$ we obtain that $P(f_{k})$ for all $k\in\mathbb{N}$. But $P$ is inclusive, and therefore ${\mathrm{f}ix}(T)\in P$.
7.4 Shrinking Maps
If $(M,d)$ is a spherically complete ultrametric space, then every shrinking map $T:M\to M$ has a unique fixpoint ([43], Theorem 1). The proof of this result relies on Zorn’s lemma for showing the existence of a maximal, with respect to set inclusion, ball $B_{z}$ in the set of balls of the form $B_{x}\,\coloneqq\,\mathrm{B}[{x},{d(x,T(x))}]$ for $x\in X$; this $z\in X$ is the unique fixpoint of $T$. A ”more constructive proof”, not relying on Zorn’s lemma, also shows that there is fixpoint of $T$ in every ball of the form $\mathrm{B}[{x},{d(x,T(x))}]$ for $x\in X$ (Corollary 5 in [33], see also [34], Chapter 5.5).
The following statement follows directly from Theorem 2.1 in [35] and the fact that ${{A}^{\omega,{n}}}$ is a spherically complete ultrametric space 2. This extension of the results in [43] for multivalued functions also relies on the application of Zorn’s lemma. Notice that these results are stated for complete non-Archimedean normed spaces, but they evidently hold also in case of spherically complete ultrametric spaces.
Lemma 14
$T:{{A}^{\omega,{n}}}\to{\mathrm{Comp}({{{A}^{\omega,{n}}}})}$ has a fixpoint if
$\mathcal{H}_{d}(T(f),T(g))$ $<d(f,g)$ (30)
@@ 7.4 Shrinking Maps
Since every singleton set is compact, we obtain the existence of a fixpoint, in particular, for deterministic maps. Uniqueness of this fixpoint follows, for example, from ([33]; see also [34], Theorem 5.4).
Corollary 7
Every strongly causal stream transformer $T:{{A}^{\omega,{n}}}\to{{A}^{\omega,{n}}}$ has a unique fixpoint.
Remark 23
The fixpoint for shrinking deterministic stream transformers in Corollary 7 is obtained as the limit of a transfinite iteration ([50]; see also [34], Remark 5.5). Now, one obtains an induction principle for shrinking (and therefore also strongly causal) deterministic stream transformers analogously to Lemma 13.
7.5 Non-expansive Maps
In the light of Lemma 7 we can use causality instead of the equivalent non-expansion property ($\mathcal{H}_{d}(T(f),T(g))\leq d(f,g)$, for all $f,g$) in Theorem 2.2 of [35].
Theorem 7.4
If $T:{{A}^{\omega,{n}}}\to{\mathrm{Comp}({{{A}^{\omega,{n}}}})}$ is causal then either $T$ has a fixpoint or there exists a ball $B$ with radius $r>0$ that that $d(u,T(u))=r$ for all $u\in B$.
Example 25
@@ 7.5 Non-expansive Maps
Theorem 7.4 provides a concise verification condition for establishing the existence of fixpoint for causal stream transformers in SAL such as the one discussed in Example 23.
8 Remarks
Stream transformers may be specified using a rich set of interrelated formalisms, including
  1. 1.
    contractual stream transformers,
  2. 2.
    rational stream transformers,
  3. 3.
    formal power series as transfer functions,
  4. 4.
    co-inductive definitions,
  5. 5.
    behavioral differential equations,
  6. 6.
    Mealy machines, and
  7. 7.
    finite stream circuits.
For instance, for every causal stream transformer one may construct a corresponding (not necessarily finite-state) Mealy machine, and vice versa. There is a similar correspondence between rational stream transformers and finite stream cirucits. In this way, there is a duality between the functional and a more machine-oriented view of causal stream transformers. This duality may advantageously be used in system design for switching between these intertwined view points as needed.
Based on the equivalence of causality and contraction for vector- and multivalued maps we have been formulating fixpoint results for nondeterministic stream transformers. Hereby, it is key that the underlying space of streams is equipped with the prefix ultrametric, which is spherically complete. In this way we obtain fixpoints for strongly causal, and therefore contractive, stream transformers with compact codomains (Theorem 7.2). If such a multivalued stream transformer is, in addition, also strongly contractive then the fixpoint is unique. Also notice that the definitions of weak and strong contraction for multivalued maps, respectively, resemble powerdomain constructions as commonly used for modeling nondeterminism in the denotational semantics of programming languages [1].
@@ 8 Remarks
The statement and proof of Theorem 7.2 is based on similar fixpoint result for multivalued maps by Nadler ([38], Theorem 5). Generalizations are described, for example, in [49, 20, 24, 51]. In particular, the papers [49, 48] are concerned with ultrametric spaces whose distance functions take their values in an arbitrary partially ordered set, not just in the real numbers, and Theorem 7.2 may also have been obtained from ([49], 3.1), but here we prefer to stay in the framework of real-valued metrics.
@@
Khamsi [31] uses a generalization of the Banach fixpoint principle to multivalued maps on a Cauchy complete metric space for developing a fixpoint semantics of stratified disjunctive logic programs (see also [26]). We are not aware, however, of previous attempts for developing fundamental concepts of system design based on multivalued fixpoints principles in the ultrametric space of streams.
@@
The derivation of the fixpoint induction principle (Lemma 13) relies on fixpoint iteration. This induction principle is reminiscent, of course, of the fixpoint induction principle for Scott-continuous functions and Kleene’s fixpoint theorem [62]. Moreover, the induction principle for multi-valued stream transformers is analogous to Park’s lemma, which is an immediate consequence of the Knaster-Tarski fixpoint theorem for monotone functions on complete lattices.
We may also develop metric-based fixpoint approximations of stream transformers over a concrete domain by corresponding fixpoints in an abstracted domain as follows. A stream abstraction $\alpha:{C}^{\omega}\to{A}^{\omega}$ relates concrete $C$-valued streams with abstract $A$-valued streams, and a corresponding stream concretization $\gamma:{A}^{\omega}\to\mathcal{P}({C}^{\omega})\setminus\{\emptyset\}$ maps abstract streams to a corresponding set of concrete streams. Such a pair $(\alpha,\gamma)$ forms, for $c$ a constant, what we may call a $c$-bounded Galois connection between ${C}^{\omega}$ and ${A}^{\omega}$ if and only if
$d(\alpha(f),g)<c$  iff $\text{~{}~{}~{}iff~{}~{}~{}}d(f,\gamma(g))<c$ (31)
for all $f\in{C}^{\omega}$ and $g\in{A}^{\omega}$. Now, for strongly causal (deterministic) stream transformers $T_{C}:{C}^{\omega}\to{C}^{\omega}$ and $T_{A}:{A}^{\omega}\to{A}^{\omega}$ such that $d(T_{C},\gamma\,\circ\,T_{A}\,\circ\,\alpha)<c$ one obtains $d({\mathrm{fix}}({T_{C}}),\gamma({\mathrm{fix}}({T_{A}})))<c$ for approximating the concrete fixpoint by an abstract fixpoint. These kinds of fixpoint approximation are useful for stream verification procedures based on model checking [14, 56, 13] and symbolic execution [32, 28] for the WP (or SP) stream calculus.
@@
Vector-valued metrics with codomain $\mathbb{R}^{n}$, for $n\geq 1$, are a viable alternative to our use of the supremum valuation for products of streams. The classical Banach contraction principle was extended by Perov [42] for contraction mappings on spaces endowed with (finite dimensional) vector-valued metrics with codomain $\mathbb{R}^{n}$. Generalizations to multivalued contraction mappings have been developed, among others, by Filip and Perusel [19, 2]. These fixpoint theorems rely on contractive mappings. Therefore, they are applicable to strongly causal but not to weakly causal stream transformers. The advantage of taking the supremum valuation is that fixpoint results for multivalued mappings apply readily, whereas the use of vector-valued metrics requires explicit extensions of these results to this modified setting. On the other hand, contractions of endomorphisms with respect to vector-valued metric can be defined more generally by requiring a square matrix $L$ with nonnegative entries such that $L^{k}\to 0$ as $k\to\infty$. Here we restricted ourselves to the cases where $L$ is of the specific form $l\cdot I$ for $l<1$ a Lipschitz constants and $I$ the identity matrix.
Additional restrictions are needed for the existence of fixpoints for causal transformers. Consider, for example, the Browder-Göhde-Kirk fixpoint theorem: for $C$ a closed, convex, bounded subset of a uniformly convex Banach space, every nonexpansive $T:C\to C$ has at least one fixpoint. In fact, if $x_{0}$ is any point in $C$, and a sequence $(x_{k})_{k\in\mathbb{N}}$ is defined by $x_{k+1}=T(x_{k})$, then the asymptotic center222 For the mapping $f:C\to[0,\infty)$ such that $u\mapsto limsup_{k\to\infty}d(u,x_{k})$, the unique point $u_{0}\in C$ satisfying $f(u_{0})=min(f(C))$ is the asymptotic center of the sequence $(x_{k})$ with respect to the set $C$. of the sequence $(x_{k})$ with respect to $C$ is a fixpoint of $F$ [57]. Since $({A}^{\omega},d)$ is uniformly convex (see Appendix 0.A), the Browder-Göhde-Kirk fixpoint principle applies to non-expansive, and therefore weakly causal, stream transformers. For an overview on fixpoints for nonexpansive maps see [30, 34].
The causality condition on stream maps is rather restrictive as it includes continuity. This condition may be relaxed, however, since the space $({A}^{\omega},d)$ is $\varepsilon$-chainable (for given $\varepsilon>0$) (Appendix 0.D), which intuitively states that there is a finite path between any two streams with intermediate ”jumps” less than $\varepsilon$. For $\varepsilon$-chainable spaces a local contraction condition suffices to obtain fixpoints for single- [18] and multivalued maps [38].
Our developments may also be extended to dense, $A$-valued streams $(a_{t})_{t\in\mathbb{R}}$ by defining valuations on dense streams along the line of valuations for discrete streams. In this way, one may also mix dense and discrete streams in heterogeneous products of streams for modeling hybrid systems. Moreover, probabilistic systems are modeled by Menger’s probabilistic metric space (or any variant thereof), in which the distance between any two points is a probability distribution function [58, 36]. Such an endeavor yields a comprehensive mathematical foundation for the formal construction of cyber-physical systems [37, 9] and their realization on conventional (von Neumann, dataflow, quantum) computers, but eventually also on dedicated, still to be developed, hypercomputing [40] machinery for capturing unbounded nondeterminism.
9 Conclusions

### v2

@@ 7.1. Set Transformers
Corollary 7.8 (Induction Principle for ${\mathrm{sp}}$).
Let $P\subseteq{{A}^{\omega,{n}}}$ be a closed set of streams and $T:{{A}^{\omega,{n}}}\to{\mathrm{Comp}({{{A}^{\omega,{n}}}})}$ contracting; then
(24) $(\forall{Q\subseteq P})\,{{\mathrm{sp}}_{T}(Q)\subseteq P}$ $\,\Rightarrow\,{\mathrm{fix}}({sp_{T}})\subseteq P\mbox{\@.}$
Proof.
For $F_{0}\,\coloneqq\,\emptyset$ and $F_{k+1}={\mathrm{sp}}_{T}(F_{k})$ we get, by induction on $k$ and the assumption above, that $F_{k}\subseteq P$ for all $k\in\mathbb{N}$. The closed set $P$ contains all of its limit points, and therefore ${\mathrm{fix}}({{\mathrm{sp}}_{T}})=\lim_{k\to\infty}F_{k}\subseteq P$. ∎
For the adjointness of ${\mathrm{sp}}_{T}$ and ${\mathrm{wp}}_{T}$ we might also use the equivalent assumption $(\forall{Q\subseteq P})\,{Q\subseteq{\mathrm{wp}}_{T}(P)}$ in the induction principle for ${\mathrm{sp}}_{T}$ (see Corollary 7.8).
7.2. Contractive Maps
The constant map $T_{\emptyset}$ to the emptyset can not have fixpoints, since $f\notin\emptyset=T_{\emptyset}(f)$. Also, we restrict the codomain of multivalued mappings to closed sets only, since any closed subset of a Cauchy complete space is Cauchy complete.
Lemma 7.9.
A nondeterministic contractive stream transformer $T$ has a fixpoint if (1) $T$ is not the constant map to the empty set, and (2) $T(f)$ is closed for all streams $f$.
Proof.
Since $T$ is not the constant map to the empty set, there exist $f_{0},f_{1}\in{A}^{\omega}$ such that $f_{1}\in T(f_{0})$. Now, since $T$ is contractive there is $10\leq l<$ 1 and a $f_{2}\in T(f_{1})$ with $d(f_{1},f_{2})\leq l\cdot d(f_{0},f_{1})$. In this way, we recursively construct a sequence $(f_{k})$ such that $f_{k+1}\in T(f_{k})$ and
@@ 7.2. Contractive Maps
(25) $d(f_{k},f_{m})\leq\mathrm{max}(d(f_{k},f_{k+1}),\ldots,d(f_{m-1},f_{m}))\leq l^{k}\mbox{\@,}$
@@
(26) $d(f_{k},f^{*})$ $\leq l^{k}\mbox{\@.}$
Since $T$ is a contraction, there is a sequence $(g_{k})_{k\in\mathbb{N}}$ such that the inequality $d(f_{k+1},g_{k})\leq l\cdot d(f_{k},f^{*})$ holds. Therefore, by inequality (26):
@@
Now, $\lim_{k\to\infty}g_{k}=f^{*}$, since $\lim_{k\to\infty}d(f^{*},g_{k})=0$. But $g_{k}\in T(f^{*})$ and all images of $T$ are closed, and therefore the limit $f^{*}$ of the sequence $(g_{k})$ is an element of $T(f^{*})$. Altogether, $f^{*}\in T(f^{*})$. ∎
In general, fixpoints of contractive multivalued maps in the sense of Definition 6.1 are not unique, but a ”slightly” stronger requirement on contractiveness implies uniqueness of fixpoints.
Definition 7.10 (Strong Contractions).
@@
If $\emptyset\notin T({{A}^{\omega,{n}}})$ then every strong $l$-contraction also is a (weak) $l$-contraction in the sense of Definition 6.1. Moreover, a deterministic stream transformer $T$ is $l$-contractive if and only if it is strongly $l$-contractive.
Lemma 7.11.
@@
Thus, $d(f^{*},g^{*})=0$, and any two fixpoints of $T$ are equal. ∎
For deterministic maps, the fixpoint iteration with $f_{k+1}\in T(f_{k})$, as constructed in the proof of Lemma 7.9 reduces to the Picard iteration $f_{k+1}=T(f_{k})$. Moreover, uniqueness of fixpoints for deterministic, $\delta>0$-causal maps can also be shown directly, since the equality of two arbitrary fixpoints $f^{*}$, $g^{*}$ directly follows from $d(f^{*},g^{*})=d(T(f^{*}),T(g^{*}))\leq 2^{-\delta}\cdot d(f^{*},g^{*})$. The following result directly follows from Lemmata 6.27.9, and 7.11.
Theorem 7.12.
A strongly causal stream transformer $T$ has a fixpoint if it is (1) not the constant map to the empty set, and (2) $T(f)$ is closed for each stream $f$. If, in addition, $T$ is strongly contractive then this fixpoint is unique.
As an immediate consequence of Theorem 7.12 we obtain as a special case a reformulation of Banachs’s fixpoint principle for streams.
Corollary 7.13.
@@
Fixpoint Induction.
The fixpoint iteration for strongly causal stream transformers in Theorem 7.12 is used to derive a fixpoint induction principle. The overall approach is analogous to deriving fixpoint induction for Scott-continuous functions on complete partial orders (Winskel, 1993). for strongly causal stream transformers.
Lemma 7.14 (Fixpoint Induction).
Let $T$ be a strongly causal stream transformer as in Theorem 7.12 with a unique fixpoint ${\mathrm{f}ix}(T)$ and $P$ a closed set of streams; then:
(27) $((\forall{f\in P})\,{T(f)\in P})\,\Rightarrow\,{\mathrm{f}ix}(T)\in P\mbox{\@.}$
Proof.
Let $(f_{k})$ be the sequence with $f_{k+1}\in T(f_{k})$ for $k\in\mathbb{N}$ and ${\mathrm{f}ix}(T)=\lim_{k\to\infty}f_{k}$ as constructed in the proof of Lemma 7.9. From the assumption and natural induction on $k$ we obtain that $f_{k}\in P$ for all $k\in\mathbb{N}$. But $P$ is closed, and therefore ${\mathrm{f}ix}(T)\in P$. ∎
7.3. Shrinking Maps
If $(M,d)$ is a spherically complete ultrametric space, then every shrinking map $T:M\to M$ has a unique fixpoint ((Petalas and Vidalis, 1993), Theorem 1). The proof of this result relies on Zorn’s lemma for showing the existence of a maximal, with respect to set inclusion, ball $B_{z}$ in the set of balls of the form $B_{x}\,\coloneqq\,\mathrm{B}[{x},{d(x,T(x))}]$ for $x\in X$; this $z\in X$ is the unique fixpoint of $T$. A ”more constructive proof”, not relying on Zorn’s lemma, also shows that there is fixpoint of $T$ in every ball of the form $\mathrm{B}[{x},{d(x,T(x))}]$ for $x\in X$ (Corollary 5 in (Kirk and Shahzad, 2012), see also (Kirk and Shahzad, 2014), Chapter 5.5).
The following statement follows directly from Theorem 2.1 in (Kubiaczyk and Ali, 1996) and the fact that ${{A}^{\omega,{n}}}$ is a spherically complete ultrametric space 3.3. This extension of the results in (Petalas and Vidalis, 1993) for multivalued functions also relies on the application of Zorn’s lemma. Notice that these results are stated for complete non-Archimedean normed spaces, but they evidently hold also in case of spherically complete ultrametric spaces.
Lemma 7.15.
The nondeterministic stream transformer $T:{{A}^{\omega,{n}}}\to{\mathrm{Comp}({{{A}^{\omega,{n}}}})}$ has a fixpoint if
(28) $\mathcal{H}_{d}(T(f),T(g))$ $<d(f,g)$
@@ 7.3. Shrinking Maps
Since every singleton set is compact, we obtain the existence of a fixpoint, in particular, for deterministic maps. Uniqueness of this fixpoint follows, for example, from ((Kirk and Shahzad, 2012); see also (Kirk and Shahzad, 2014), Theorem 5.4).
Corollary 7.17.
Every strongly causal deterministic stream transformer $T:{{A}^{\omega,{n}}}\to{{A}^{\omega,{n}}}$ has a unique fixpoint.
The fixpoint for shrinking deterministic stream transformers in Corollary 7.17 is obtained as the limit of a transfinite iteration ((Prieß-Crampe and Ribenboim, 2013); see also (Kirk and Shahzad, 2014), Remark 5.5). Now, one obtains an induction principle for shrinking (and therefore also strongly causal) deterministic stream transformers analogously to Lemma 7.14.
7.4. Nonexpansive Maps
In the light of Lemma 6.9 we can use causality instead of the equivalent nonexpansion property ($\mathcal{H}_{d}(T(f),T(g))\leq d(f,g)$, for all $f,g$) in Theorem 2.2 of (Kubiaczyk and Ali, 1996).
Theorem 7.18.
If the nondeterministic stream transformer $T:{{A}^{\omega,{n}}}\to{\mathrm{Comp}({{{A}^{\omega,{n}}}})}$ is causal then either $T$ has a fixpoint or there exists a ball $B$ with radius $r>0$ that that $d(u,T(u))=r$ for all $u\in B$.
Example 7.19.
@@ 7.4. Nonexpansive Maps
Theorem 7.18 provides a concise verification condition to establish the existence of fixpoints for causal stream transformers in SAL such as the one discussed in Example 5.10.
8. Related Work
The causality-contraction correspondence for the special case of deterministic stream transformers follows from a straightforward unfolding of the definitions. This was already observed by de Bakker and Zucker (de Bakker and Zucker, 1982), as the basis for using Banach’s fixed point theorem to replace the least fixed point approach of denotational semantics based on complete partial orders. In particular, de Bakker’s notion of guardedness is closely related to strong causality (de Bakker, 2001). The correspondence between causality and contraction for deterministic stream transformers has also been heavily used in the context of functional reactive programming (for, example (Krishnaswami and Benton, 2011; Birkedal et al., 2012)) for solving recursive stream equations. Broy ((Broy, 2023), A.3) uses this correspondence to solve stream inclusions on timed streams of the form $x\in F(y)$. In his approach, stream inclusions are replaced with stronger equalities of the form $\bar{x}=g(\bar{y})$, where $g$ is a so-called strongly causal representation of $F$ and $\bar{x}$, $\bar{y}$ are untimed streams corresponding to $x$ and $y$, respectively. Unlike previous work, we extend the causality-contraction correspondence to nondeterministic stream transformers and solve causal stream inclusions directly by corresponding fixed-point results motivated by results from non-Archimedean functional analysis.
@@ 8. Related Work
The statement and proof of Theorem 7.12 is based on analogous fixpoint results for multivalued maps by Nadler ((Nadler, 1969), Theorem 5). Generalizations are described, for example, in (Prieß-Crampe and Ribenboim, 2000; Gajić, 2001; Hajimojtahed and Mirmostafaee, 2016; Rao and Kishore, 2008). In particular, the papers (Prieß-Crampe and Ribenboim, 2000, 1993) are concerned with ultrametric spaces whose distance functions take their values in an arbitrary partially ordered set, not just in the real numbers, and Theorem 7.12 may also have been obtained from the results ((Prieß-Crampe and Ribenboim, 2000), 3.1), but here we prefer to stay in the framework of real-valued metrics.
@@
Khamsi (Khamsi et al., 1993) uses a generalization of the Banach fixpoint principle to multivalued maps on a Cauchy complete metric space for developing a fixpoint semantics of stratified disjunctive logic programs (see also (Hitzler and Seda, 2010)). We are not aware, however, of previous attempts to develop fundamental concepts of system design based on multivalued fixpoints principles in the ultrametric space of streams.
@@
The derivation of the fixpoint induction principle (Lemma 7.14) relies on fixpoint iteration. This induction principle is reminiscent, of course, of the fixpoint induction principle for Scott-continuous functions and Kleene’s fixpoint theorem (Winskel, 1993). Moreover, the induction principle for multivalued stream transformers is analogous to Park’s lemma, which is an immediate consequence of the Knaster-Tarski fixpoint theorem for monotone functions on complete lattices.
@@
Vector-valued metrics with codomain $\mathbb{R}^{n}$, for $n\geq 1$, are a viable alternative to the use of the supremum valuation for products of streams. The classical Banach contraction principle was extended by Perov (Perov, 1964) for contraction mappings on spaces endowed with (finite dimensional) vector-valued metrics with codomain $\mathbb{R}^{n}$. Generalizations to multivalued contraction mappings have been developed, among others, by Filip and Perusel (Filip and Petruşel, 2010; Almalki et al., 2022). These fixpoint theorems rely on contractive mappings. Therefore, they are applicable to strongly causal but not to weakly causal stream transformers. The advantage of taking the supremum valuation is that fixpoint results for multivalued mappings apply readily, whereas the use of vector-valued metrics requires explicit extensions of these results to this modified setting. On the other hand, contractions of endomorphisms with respect to vector-valued metric can be defined more generally by requiring a square matrix $L$ with nonnegative entries such that $\lim_{k\to\infty}L^{k}=0$. Here we restricted developments to the cases where $L$ is of the specific form $l\cdot I$ for $l<1$ a Lipschitz constants and $I$ the identity matrix.
More constructive proofs of fixpoints of deterministic transformers on spherically complete ultrametric spaces result in transfinite Picard iterations (Khamsi and Kirk, 2011; Kirk and Shahzad, 2014). These kinds of results could therefore be used to formulate induction principles for the solutions of deterministic causal stream transformers. But we are not aware of general results in this direction for nondeterministic maps.
9. Conclusions